OpenStudy (anonymous):

Solve the following simultaneous equations: 3x-2y=y(x+2y)=4 @ganeshie8 @perl

OpenStudy (perl):

this is one equation? 3x-2y=y(x+2y)=4

OpenStudy (anonymous):

The question written 3x-2y=y(x+2y)=4

OpenStudy (anonymous):

just give me a few minutes :)

OpenStudy (perl):

sure

OpenStudy (perl):

ok i see it

OpenStudy (anonymous):

Any ideas?

OpenStudy (perl):

lets break apart the equality 3x-2y=y(x+2y)=4 becomes 3x-2y = 4 y(x+2y) = 4

OpenStudy (anonymous):

okay

OpenStudy (perl):

Lets label the equations, equation 1 and equation 2 1) 3x-2y = 4 2) y(x+2y) = 4 In the first equation solve for x 3x - 2y = 4 x = ( 4 + 2y) / 3 In the second equation distribute or expand yx + 2y^2 = 4

OpenStudy (perl):

ok :)

OpenStudy (anonymous):

Thnx @perl :)